# On discrete Sobolev inequalities for nonconforming finite elements under a semi-regular mesh condition

### v1

@@ 2.3 Finite element spaces
Let $Ne$ be the number of elements included in the mesh $\mathbb{T}_{h}$. Thus, we write $\mathbb{T}_{h}=\{T_{j}\}_{j=1}^{Ne}$. We introduce a discontinuous CR finite element space.
@@
$\displaystyle V_{h}^{DCCR}$ $:=\left\{\sum_{j=1}^{Ne}\sum_{i=1}^{d+1}c_{j(i)}\phi_{j(i)};\ c_{j(i)}\in\mathbb{R},\ \forall i,j\right\}\subset V_{h}^{DC(1)}.$ (2.6)
@@ 2.4 Norms
Let $F\in\mathcal{F}_{h}^{i}$ with $F=T_{+}\cap T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$, +>+>- be an interior face and $F\in\mathcal{F}_{h}^{\partial}$ with $F=\partial T_{\partial}\cap\partial\Omega$, $T_{\partial}\in\mathbb{T}_{h}$ a boundary face. Let $p\in[1,\infty)$. A choice for the weighted parameters is such that
@@
where $\ell_{T,F}$ is defined in Section 2.2. See [7, Section 2.4] for more information on this parameter.
Let $p\in[1,\infty)$ and $V_{h}\in\{V_{h}^{DCCR},V_{h}^{CR},V_{h0}^{CR}\}$. We define the following three norms for any $\varphi_{h}\in V_{h}$;
@@
We note that
$\displaystyle|\varphi_{h}|_{p,J}$ $=\left(\sum_{F\in\mathcal{F}_{h}^{\partial}}\kappa_{p,F*}\|\Pi_{F}^{0}[\![\varphi_{h}]\!]\|_{L^{p}(F)}^{p}\right)^{\frac{1}{p}}\quad\text{if $\varphi_{h}\in V_{h}^{CR}$},$
$|\varphi_{h}|_{p,J}$ $=0\quad\text{if $\varphi_{h}\in V_{h0}^{CR}$}.$
@@ 2.5.2 Three-dimensional case
In the three-dimensional case, we consider the following two cases: (i) and (ii); see Condition 2.4.
@@ 2.6 Two-step affine mapping
In anisotropic meshes, the mesh shape and element proportions are non-uniform, which directly affects the interpolation accuracy. Existing interpolation error estimates typically assume an even or regular mesh, which may overestimate the error if applied directly to anisotropic meshes. To remedy this, we proposed in [6, 8, 13, 15] a new strategy on anisotropic meshes.
@@ 2.6.2 Construct mapping $\Phi_{T}:\widetilde{T}\to T$
Condition 2.3 (Case in which $d=2$).
@@
Condition 2.4 (Case in which $d=3$).
@@
A proof can be found in [15, Lemma 2]. ∎
@@ 2.9 $L^{2}$-orthogonal projection
This section considers error estimates of the $L^{2}$-orthogonal projection, e.g., for a standard argument, see [4, Section 11.5.3]. Here, the discussion is based on the two-step affine mapping.
@@
The $L^{2}$-orthognal projections onto $\widehat{P}$ and $P$ are respectively the linear operators $\Pi^{0}_{\widetilde{T}}:L^{1}(\widetilde{T})\to\widetilde{P}$ and $\Pi^{0}_{{T}}:L^{1}({T})\to{P}$ defined as
@@
Because all the norms in the finite-dimensional space $\widehat{P}$ are equivalent, there exist $\hat{c}_{1}$ and $\hat{c}_{2}$, depending on $\widehat{T}$, such that
$\displaystyle\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}$ $\leq\hat{c}_{1}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{2}(\widehat{T})}$, (2.22)
$\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q^{\prime}}(\widehat{T})}$ $\leq\hat{c}_{2}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}$, (2.23)
where $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. Then,
$\displaystyle\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}^{2}$ $\leq c\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{2}(\widehat{T})}^{2}=c\int_{\widehat{T}}\hat{\varphi}\Pi_{\widehat{T}}^{0}\hat{\varphi}d\hat{x}$
$\leq c\|\hat{\varphi}\|_{L^{q}(\widehat{T})}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q^{\prime}}(\widehat{T})}$
$\leq c\|\hat{\varphi}\|_{L^{q}(\widehat{T})}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}$,
where we used (2.22), (2.23), (2.20) with $\hat{q}:=\Pi_{\widehat{T}}^{0}\hat{\varphi}$, and the Hölder’s inequality with $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. This proves the target inequality. ∎
@@
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}\leq c|T|_{d}^{\frac{1}{q}-\frac{1}{p}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\varphi}{\partial\bm{r}_{i}}\right\|_{L^{p}(T)}.$ (2.25)
@@
Using the scaling argument, we have
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}$ $=c|\det({A}_{\widetilde{T}})|^{\frac{1}{q}}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}-\hat{\varphi}\|_{L^{q}(\widehat{T})}$. (2.26)
where we used $|\det(A_{T})|=1$. For any $\hat{\eta}\in\mathbb{P}^{0}\subset\widehat{P}$, from the triangle inequality and $\Pi_{\widehat{T}}^{0}\hat{\eta}=\hat{\eta}$, we have
@@
$\displaystyle\|\Pi_{\widehat{T}}^{0}(\hat{\varphi}-\hat{\eta})\|_{L^{q}(\widehat{T})}\leq c\|\hat{\varphi}-\hat{\eta}\|_{L^{q}(\widehat{T})}$. (2.28)
Using the Sobolev embedding theorem for the second term on the right-hand side of (2.27) and (2.28), we obtain
$\displaystyle\|\hat{\varphi}-\hat{\eta}\|_{L^{q}(\widehat{T})}\leq c\|\hat{\varphi}-\hat{\eta}\|_{W^{1,p}(\widehat{T})}$. (2.29)
@@
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}$ $\leq c|\det({A}_{\widetilde{T}})|^{\frac{1}{q}}\inf_{\hat{\eta}\in\mathbb{P}^{0}}\|\hat{\varphi}-\hat{\eta}\|_{W^{1,p}(\widehat{T})}$. (2.30)
[…]
$\displaystyle|\hat{\varphi}-\hat{\eta}_{\beta}|_{W^{t,p}(\widehat{T})}\leq C^{BH}(\widehat{T})|\hat{\varphi}|_{W^{1,p}(\widehat{T})},\quad t=0,1.$ (2.31)
Using the new scaling argument [15, Lemma 6] ($\ell=1$ and $m=0$) and (2.31), we then have
$\displaystyle\|\hat{\varphi}-\hat{\eta}_{\beta}\|_{W^{1,p}(\widehat{T})}$ $\leq c|\hat{\varphi}|_{W^{1,p}(\widehat{T})}$
$\leq c|\det({A}_{\widetilde{T}})|^{-\frac{1}{p}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\varphi}{\partial\bm{r}_{i}}\right\|_{L^{p}(T)}.$ (2.32)
Therefore, combining (2.30) and (2.32), we have (2.25). Here, we used
@@ 3.2 RT finite element interpolation operator
where ${\bm{\theta}}_{T,i}^{RT}$ is the local shape basis function (e.g. [4, p. 162]) and ${\bm{n}}_{T,i}$ is a fixed unit normal to ${F}_{T,i}$.
@@
$\displaystyle\|{I_{T}^{RT}\bm{v}}\|_{L^{p}({T})^{d}}$ $\leq c\left[\frac{H_{T}}{h_{T}}\left(\|\bm{v}\|_{L^{p}(T)^{d}}+\sum_{|\varepsilon|=1}h^{\varepsilon}\left\|\partial_{\bm{r}}^{\varepsilon}\bm{v}\right\|_{L^{p}(T)^{d}}\right)+h_{T}\|\bm{\nabla}\cdot\bm{v}\|_{L^{p}({T})}\right].$ (3.4)
@@
A proof can be found in [6, Lemma 8]. ∎
@@
$\displaystyle\|{I_{T}^{RT}\bm{v}}\|_{L^{p}({T})^{3}}$ $\leq c\frac{H_{T}}{h_{T}}\left[\|\bm{v}\|_{L^{p}(T)^{3}}+h_{T}\sum_{k=1}^{3}\left\|\frac{\partial\bm{v}}{\partial\bm{r}_{k}}\right\|_{L^{p}(T)^{3}}\right].$ (3.5)
@@
A proof can be found in [6, Lemma 8]. ∎
@@
Let $p\in[1,\infty)$. We impose Assumption 1 with $h\leq 1$. Then,
$\displaystyle\|{I_{h}^{RT}\bm{v}}\|_{L^{p}({\Omega})^{d}}$ $\leq C(\gamma_{0})\|\bm{v}\|_{W^{1,p}(\Omega)^{d}}\quad\forall\bm{v}\in W^{1,p}(\Omega)^{d}.$
@@
$\displaystyle\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{d}}$ $\leq c\left(\frac{H_{T}}{h_{T}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\bm{v}}{\partial r_{i}}\right\|_{L^{p}(T)^{d}}+h_{T}\|\mathop{\mathrm{div}}\bm{v}\|_{L^{p}({T})}\right).$ (3.6)
@@
A proof can be found in [6, Theorem 2]. ∎
@@
$\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{3}}\leq c\frac{H_{T}}{h_{T}}\Biggl{(}h_{T}|\bm{v}|_{W^{1,p}(T)^{3}}\Biggr{)}.$ (3.7)
@@
A proof can be found in [6, Theorem 3]. ∎
@@
$\displaystyle\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{d}}$ $\leq C_{RT}(\gamma_{0})h_{T}|\bm{v}|_{W^{1,p}(T)^{d}}.$ (3.8)

### v2

@@ 2.3 Finite element spaces
In this paper, we focus exclusively on the lowest-order cases, specifically $k=0$ and $k=1$. The space $V_{h}^{\mathrm{DC}(0)}$, consisting of piecewise constant functions, is utilised through the elementwise $L^{2}$-projection $\Pi_{h}^{0}:L^{1}(\Omega)\to V_{h}^{\mathrm{DC}(0)}$. Meanwhile, the CR spaces $V_{h}^{\mathrm{CR}}$, $V_{h0}^{\mathrm{CR}}$, and the discontinuous CR space $V_{h}^{\mathrm{DCCR}}$, which are introduced subsequently, are subspaces of $V_{h}^{\mathrm{DC}(1)}$. The discrete $L^{q}-L^{p}$ Sobolev inequalities demonstrated in Section 5 are applicable solely to these lowest-order spaces. The limitation to $k=0,1$ is not solely a matter of notation. Extending the current framework to encompass higher-order CR-type elements ($k\geq 2$) on semi-regular anisotropic meshes necessitates the development of a corresponding higher-order CR-type theory. This would include anisotropic approximation and interpolation error estimates based on the parameters of the semi-regular mesh. As such results are presently unavailable within the semi-regular anisotropic context under consideration, we do not pursue this extension.
Recalling that $\mathbb{T}_{h}=\{T_{j}\}_{j=1}^{Ne}$, we introduce a discontinuous CR finite element space.
@@
$\displaystyle V_{h}^{DCCR}$ $:=\left\{\sum_{j=1}^{Ne}\sum_{i=1}^{d+1}c_{j(i)}\phi_{j(i)};\ c_{j(i)}\in\mathbb{R},\ \forall i,j\right\}{=}V_{h}^{DC(1)}.$ (2.6)
@@
Remark 2.4.
On each element $T_{j}$, the local CR basis functions $\{\theta_{T_{j},i}^{CR}\}_{i=1}^{d+1}$ form a basis of $\mathbb{P}^{1}({T}_{j})$, and $\phi_{j(i)}$ are exactly these functions extended by zero outside $T_{j}$. Therefore, $V_{h}^{DCCR}=V_{h}^{DC(1)}$. We keep the notation $V_{h}^{DCCR}$ to emphasise that we use CR-type degrees of freedom and basis functions on each element.
@@ 2.4 Norms
Let $F\in\mathcal{F}_{h}^{i}$ with $F=\partial T_{+}\cap\partial T_{-}$, $T_{+},T_{-}\in\mathbb{T}_{h}$, +>+>- be an interior face and $F\in\mathcal{F}_{h}^{\partial}$ with $F=\partial T_{\partial}\cap\partial\Omega$, $T_{\partial}\in\mathbb{T}_{h}$ a boundary face. Let $p\in[1,\infty)$. A choice for the weighted parameters is such that
@@
The weights $\omega_{T_{i},F}$ satisfy $\omega_{T_{i},F}\in[0,1]$ and $\omega_{T_{+},F}+\omega_{T_{-},F}=1$ and they depend only on the face heights $\ell_{T_{i},F}$. They will be used in the definition of weighted averages on interior faces so that the contribution of each neighbouring element is tuned according to its height in the normal direction to $F$.
@@
where $\ell_{T,F}$ is the height of $T$ with respect to $F$ in Section 2.2. This parameter is constructed solely from the local face heights $\ell_{T,F}$ and plays the role of an inverse normal mesh size on each face. On a regular (shape-regular and quasi-uniform) mesh, one has $\kappa_{p,F}^{*}\approx h_{F}^{-1}\approx h^{-1}$, so that $\kappa_{p,F}^{*}$ grows as the mesh is refined. For an explicit formula and several examples illustrating its behaviour on anisotropic meshes, we refer to [13, Section 2.4].
Let $p\in[1,\infty)$ and $V_{h}\in\{V_{h}^{DCCR},V_{h}^{CR},V_{h0}^{CR}\}$. We define the following mesh-dependent functionals for any $\varphi_{h}\in V_{h}$;
@@
Remark 2.5.
The definitions of $|\cdot|_{p,J}$ and $|\cdot|_{p,V_{h}}$ are applicable to any function $v\in W^{1,p}(\mathbb{T}_{h})$, thereby establishing mesh-dependent seminorms on the broken Sobolev space $W^{1,p}(\mathbb{T}_{h})$. In this study, we focus on utilising these seminorms within the finite element spaces $V_{h}\in\{V_{h}^{{DCCR}},V_{h}^{{CR}},V_{h0}^{{CR}}\}$.
In addition, within each of these discrete spaces, $|\cdot|_{p,V_{h}}$ serves as a norm. Specifically, if $|v_{h}|_{p,V_{h}}=0$, it follows that $|v_{h}|_{W^{1,p}(\mathbb{T}_{h})}=0$, indicating that $v_{h}$ is piecewise constant over $\mathbb{T}_{h}$. For the case where $V_{h}=V_{h}^{{DCCR}}$, the condition $|v_{h}|_{p,J}=0$ ensures that the constants on each element are uniform across adjacent elements and vanish on boundary faces, resulting in $v_{h}\equiv 0$. For $V_{h}=V_{h}^{{CR}}$ or $V_{h}=V_{h0}^{{CR}}$, the CR continuity conditions across interior faces, and in the latter case, the boundary moment conditions, necessitate that any piecewise constant function must be identically zero. Thus, $|\cdot|_{p,V_{h}}$ defines a norm on each of the spaces $V_{h}^{{DCCR}},V_{h}^{{CR}},V_{h0}^{{CR}}$.
@@ 2.5.2 Three-dimensional case
In the three-dimensional case, we consider the following two cases: (i) and (ii); see Condition 2.7.
@@ 2.6 Two-step affine mapping
In anisotropic meshes, the mesh shape and element proportions are non-uniform, which directly affects the interpolation accuracy. Existing interpolation error estimates typically assume an even or regular mesh, which may overestimate the error if applied directly to anisotropic meshes. To remedy this, we proposed in [12, 19, 21] a new strategy on anisotropic meshes.
@@ 2.6.2 Construct mapping $\Phi_{T}:\widetilde{T}\to T$
Condition 2.6 (Case in which $d=2$).
@@
Condition 2.7 (Case in which $d=3$).
@@
A proof can be found in [21, Lemma 2].
@@ 2.9 $L^{2}$-orthogonal projection
This section considers error estimates of the $L^{2}$-orthogonal projection, e.g., for a standard argument, see [9, Section 11.5.3]. Here, the discussion is based on the two-step affine mapping.
@@
The $L^{2}$-orthogonal projections onto $\widehat{P}$ and $P$ are respectively the linear operators $\Pi^{0}_{\widetilde{T}}:L^{1}(\widetilde{T})\to\widetilde{P}$ and $\Pi^{0}_{{T}}:L^{1}({T})\to{P}$ defined as
@@
[…]
Using the Hölder’s inequality with $\frac{1}{q}+\frac{1}{q^{\prime}}=1$ yields
$\displaystyle|\Pi_{\widehat{T}}^{0}\hat{\varphi}|=\frac{1}{|\widehat{T}|_{d}}\left|\int_{\widehat{T}}\hat{\varphi}d\hat{x}\right|\leq\frac{1}{|\widehat{T}|_{d}}\|\hat{\varphi}\|_{L^{q}(\widehat{T})}\|1\|_{L^{q^{\prime}}(\widehat{T})}=\frac{1}{|\widehat{T}|_{d}}\|\hat{\varphi}\|_{L^{q}(\widehat{T})}|\widehat{T}|_{d}^{\frac{1}{q^{\prime}}}=|\widehat{T}|_{d}^{\frac{1}{q^{\prime}}-1}\|\hat{\varphi}\|_{L^{q}(\widehat{T})}$,
which leads to
$\displaystyle\|\Pi_{\widehat{T}}^{0}\hat{\varphi}\|_{L^{q}(\widehat{T})}$ $=\left(\int_{\widehat{T}}|\Pi_{\widehat{T}}^{0}\hat{\varphi}|^{q}d\hat{x}\right)^{\frac{1}{q}}=|\Pi_{\widehat{T}}^{0}\hat{\varphi}||\widehat{T}|_{d}^{\frac{1}{q}}$
$\leq|\widehat{T}|_{d}^{\frac{1}{q}+\frac{1}{q^{\prime}}-1}\|\hat{\varphi}\|_{L^{q}(\widehat{T})}=\|\hat{\varphi}\|_{L^{q}(\widehat{T})}$.
@@
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}\leq{C^{L^{2}}(d,p,q)}|T|_{d}^{\frac{1}{q}-\frac{1}{p}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\varphi}{\partial\bm{r}_{i}}\right\|_{L^{p}(T)},$ (2.23)
where $C^{L^{2}}(d,p,q)$ depends only on the dimension $d$, the exponents $p$ and $q$ and is independent of the physical element $T$, the mesh size $h$, and the semi-regular parameter.
@@
Let $p\in[1,\infty)$ and $q\in[1,\infty)$ be such that (2.22) holds. Let ${\varphi}=\hat{\varphi}\circ{\Phi}^{-1}$.
By the invariance of $P$ under the $L^{2}$-projection $\Pi_{T}^{0}$ (see Section 2.9) and by the definition of the affine mapping $\Phi:\widehat{T}\to T$, we have
$\displaystyle\Pi_{T}^{0}\varphi-\varphi=(\Pi_{\widehat{T}}^{0}\hat{\varphi}-\hat{\varphi})\circ{\Phi}^{-1},$
[…]
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}$ $=|\det({A}_{\widetilde{T}})|^{\frac{1}{q}}\|\Pi_{\widehat{T}}^{0}\hat{\varphi}-\hat{\varphi}\|_{L^{q}(\widehat{T})}$. (2.24)
where we used $|\det(A_{T})|=1$. We set
[…]
From the triangle inequality and $\Pi_{\widehat{T}}^{0}\hat{\eta}=\hat{\eta}$, we have
@@
$\displaystyle\|\Pi_{\widehat{T}}^{0}(\hat{\varphi}-\hat{\eta})\|_{L^{q}(\widehat{T})}\leq\|\hat{\varphi}-\hat{\eta}\|_{L^{q}(\widehat{T})}$. (2.26)
For the second term on the right-hand side of (2.25) and (2.26), by the Sobolev embedding theorem on the fixed reference simplex $\widehat{T}$, there exists a constant $C^{Sob}=C^{Sob}(d,p,q)>0$ such that
$\displaystyle\|\hat{\varphi}-\hat{\eta}\|_{L^{q}(\widehat{T})}\leq C^{Sob}\|\hat{\varphi}-\hat{\eta}\|_{W^{1,p}(\widehat{T})}$, (2.27)
where $C^{Sob}$ depends only on the dimension $d$, the exponents $p$ and $q$ (and on the shape of the fixed reference simplex $\widehat{T}$), and is independent of the physical element $T$, the mesh size $h$, and the semi-regular parameter.
@@
$\displaystyle\|\Pi_{T}^{0}\varphi-\varphi\|_{L^{q}(T)}$ $\leq 2C^{Sob}|\det({A}_{\widetilde{T}})|^{\frac{1}{q}}\|\hat{\varphi}-\hat{\eta}\|_{W^{1,p}(\widehat{T})}$. (2.28)
From the Poincaré–Steklov inequality (e.g., see [9, Lemma 3.24]),
$\displaystyle\|\hat{\varphi}-\hat{\eta}\|_{W^{1,p}(\widehat{T})}^{p}$ $=|\hat{\varphi}|_{W^{1,p}(\widehat{T})}^{p}+\|\hat{\varphi}-\hat{\eta}\|_{L^{p}(\widehat{T})}^{p}$
$\leq|\hat{\varphi}|_{W^{1,p}(\widehat{T})}^{p}+C^{PS}(p)h_{\widehat{T}}^{p}|\hat{\varphi}|_{W^{1,p}(\widehat{T})}^{p},$
which leads to
$\displaystyle\|\hat{\varphi}-\hat{\eta}\|_{W^{1,p}(\widehat{T})}$ $\leq C^{PS1}(d,p)|\hat{\varphi}|_{W^{1,p}(\widehat{T})}$, (2.29)
where we set
$\displaystyle C^{PS1}(d,p):=\left(1+C^{PS}(p)h_{\widehat{T}}^{p}\right)^{\frac{1}{p}}.$
Because the reference simplex $\widehat{T}$ is fixed once and for all, $h_{\widehat{T}}$ is a dimension-dependent constant and $C^{PS1}(d,p)$ depends only on $d$ and $p$.
Using the new scaling argument [21, Lemma 6] ($\ell=1$ and $m=0$), we then have
$\displaystyle|\hat{\varphi}|_{W^{1,p}(\widehat{T})}$ $\leq C^{Ish}(d,p)|\det({A}_{\widetilde{T}})|^{-\frac{1}{p}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\varphi}{\partial\bm{r}_{i}}\right\|_{L^{p}(T)},$ (2.30)
where $C^{Ish}(d,p)$ depends only on the exponent $d$ and the exponent $p$, and is independent of the physical element $T$, the mesh size $h$, and the semi-regular parameter.
Therefore, combining (2.28), (2.29) and (2.30) proves (2.23) with
$\displaystyle C^{L^{2}}(d,p,q):=2C^{Sob}(d,p,q)\cdot C^{PS}(d,p)\cdot C^{Ish}(d,p)\cdot(d!)^{\frac{1}{q}-\frac{1}{p}},$
which depends only on $d,p,q$ and is independent of the physical element $T$, the mesh size $h$, and the semi-regular parameter. Here, we used
@@ 3.2 RT finite element interpolation operator
where ${\bm{\theta}}_{T,i}^{RT}$ is the local shape basis function (e.g. [9, p. 162]) and ${\bm{n}}_{T,i}$ is a fixed unit normal to ${F}_{T,i}$.
@@
$\displaystyle\|{I_{T}^{RT}\bm{v}}\|_{L^{p}({T})^{d}}$ $\leq{C^{LRTsta1}(d,p)}\left[\frac{H_{T}}{h_{T}}\left(\|\bm{v}\|_{L^{p}(T)^{d}}+\sum_{|\varepsilon|=1}h^{\varepsilon}\left\|\partial_{\bm{r}}^{\varepsilon}\bm{v}\right\|_{L^{p}(T)^{d}}\right)+h_{T}\|\bm{\nabla}\cdot\bm{v}\|_{L^{p}({T})}\right].$ (3.4)
Here, the constant $C^{LRTsta1}(d,p)$ depends only on $d$ and $p$ and is independent of $h$ and of the aspect ratios and interior angles of the simplices.
@@
A proof can be found in [12, Lemma 8].
@@
$\displaystyle\|{I_{T}^{RT}\bm{v}}\|_{L^{p}({T})^{3}}$ $\leq{C^{LRTsta2}(d,p)}\frac{H_{T}}{h_{T}}\left[\|\bm{v}\|_{L^{p}(T)^{3}}+h_{T}\sum_{k=1}^{3}\left\|\frac{\partial\bm{v}}{\partial\bm{r}_{k}}\right\|_{L^{p}(T)^{3}}\right].$ (3.5)
Here, the constant $C^{LRTsta2}(d,p)$ depends only on $d$ and $p$ and is independent of $h$ and of the aspect ratios and interior angles of the simplices.
@@
A proof can be found in [12, Lemma 8].
@@
Let $p\in[1,\infty)$. We impose Assumption 1. Then,
$\displaystyle\|{I_{h}^{RT}\bm{v}}\|_{L^{p}({\Omega})^{d}}$ $\leq{C^{RTsta}(d,p,\gamma_{0})}\|\bm{v}\|_{W^{1,p}(\Omega)^{d}}\quad\forall\bm{v}\in W^{1,p}(\Omega)^{d}.$
Here, the constant $C^{RTsta}(d,p,\gamma_{0})$ depends only on $d$, $p$ and the semi-regular mesh parameter $\gamma_{0}$, and is independent of $h$ and of the aspect ratios and interior angles of the simplices.
@@
$\displaystyle\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{d}}$ $\leq{C^{LRTint1}(d,p)}\left(\frac{H_{T}}{h_{T}}\sum_{i=1}^{d}h_{i}\left\|\frac{\partial\bm{v}}{\partial r_{i}}\right\|_{L^{p}(T)^{d}}+h_{T}\|\mathop{\mathrm{div}}\bm{v}\|_{L^{p}({T})}\right).$ (3.6)
Here, the constant $C^{LRTint1}(d,p)$ depends only on $d$ and $p$ and is independent of $h$ and of the aspect ratios and interior angles of the simplices.
@@
A proof can be found in [12, Theorem 2].
@@
$\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{3}}\leq{C^{LRTint2}(d,p)}\frac{H_{T}}{h_{T}}\Biggl(h_{T}|\bm{v}|_{W^{1,p}(T)^{3}}\Biggr).$ (3.7)
Here, the constant $C^{LRTint2}(d,p)$ depends only on $d$ and $p$ and is independent of $h$ and of the aspect ratios and interior angles of the simplices.
@@
A proof can be found in [12, Theorem 3].
@@
$\displaystyle\|{I}_{T}^{RT}\bm{v}-\bm{v}\|_{L^{p}(T)^{d}}$ $\leq{C^{RTint}(d,p,\gamma_{0})}h_{T}|\bm{v}|_{W^{1,p}(T)^{d}}.$ (3.8)
Here, the constant $C^{{RTint}}(d,p,\gamma_{0})$ depends only on $d$, $p$ and the semi-regular mesh parameter $\gamma_{0}$, and is independent of $h$ and of the aspect ratios and interior angles of the simplices.
Remark 3.8.
In the classical shape-regular framework, the local Raviart-Thomas (RT) interpolation operator $I_{T}^{{RT}}$, as defined in (3.3), can be extended by continuity from $W^{1,1}(T)^{d}$ to the broader space
$\displaystyle W(T):=\{\bm{v}\in L^{q}(T)^{d};\ \mathop{\mathrm{div}}\bm{v}\in L^{2}(T)\}=H(\mathop{\mathrm{div}},T)\cap L^{q}(T)^{d},\quad q>2,$
as demonstrated in [5, Section 2.5.1] and [9, Section 17.2]. Specifically, for $\bm{v}\in W(T)$, all the degrees of freedom in (3.3) are well-defined, and $I_{T}^{{RT}}\bm{v}\in\mathbb{RT}^{0}(T)$.
In this paper, we operate under the semi-regular anisotropic mesh condition and employ the local estimates (3.4) and (3.5) from [12]. These bounds are established within a $W^{1,p}$-framework and incorporate directional derivatives of $\bm{v}$ in addition to its divergence, along with the two-step affine mapping and anisotropic trace inequalities. Currently, we are unable to extend such anisotropic estimates to rougher data $\bm{v}\in W(T)$ (or even to $\bm{v}\in H(\mathop{\mathrm{div}},T)$), particularly in terms of eliminating the gradient terms on the right-hand side of (3.4) while maintaining the same dependence of the constants on the semi-regularity parameter.
Consequently, throughout Section 3.2, we retain the regularity assumption $\bm{v}\in W^{1,p}(T)^{d}$ in all stability and RT interpolation estimates, including Lemmata 3.2-3.3, Corollaries 3.4-3.7, and Theorems 3.5-3.6. It is also noteworthy that our discrete $L^{q}-L^{p}$ Sobolev inequalities in Section 5 rely solely on the global stability estimate in Corollary 3.4.
We also refer to our previous work [12] for comprehensive discussions on anisotropic RT interpolation on semi-regular meshes and for counterexamples illustrating that Babuška–Aziz type techniques and purely $H(\mathop{\mathrm{div}})$-based approaches are not applicable in this context; see particularly in [12, Section 4].
Finally, we recall that our local RT interpolation estimates in Theorems 3.5-3.6 are based on the anisotropic results of [12]. In the Type (i) configuration, [12, Theorem 2] yields the genuinely anisotropic bound (3.6), expressed in terms of directional derivatives and the divergence of $v$. In the Type (ii) configuration, however, [12, Theorem 3] currently provides only an estimate of the form (3.7), involving the full $W^{1,p}(T)^{3}$-seminorm. We do not claim any refinement of this result here, and obtaining a sharper anisotropic estimate analogous to (3.6) for Type (ii) elements under the same semi-regular mesh condition appears to require additional ideas and is left open.
Remark 3.9 (On the dependence of constants).
Throughout this paper, we denote the parameter dependence of generic constants by expressions such as $C^{LRTsta1}(d,p)$ or $C^{dS0}(d,\Omega,p,q,\gamma_{0})$. Notably, these constants are independent of the mesh size $h$, as well as the element aspect ratios and interior angles. When applicable, their dependence on anisotropy is solely through the semi-regularity parameter $\gamma_{0}$.
In the context of anisotropic RT interpolation estimates, as exemplified in Lemma 3.2, we document only the dependence on pertinent parameters, without explicitly tracking all intermediate constants. While it is theoretically feasible to derive a closed-form expression in terms of $d$ and $p$, such an endeavour would necessitate extensive bookkeeping through the two-step affine/Piola transformations and auxiliary inequalities, which exceeds the scope of this paper. We intend to address explicit constant tracking in future research.